OpenStudy (anonymous):

given a triangle with the side b=60m, angleB=51.3deg. and side a= 76.45m, find angle C? a) 37.78deg. b) 97.43deg. c) 99.44deg. d) 44.8deg.

OpenStudy (anonymous):

use the cosin theorem

OpenStudy (stormfire1):

Use the Law of Sines: \[\frac{\sin(51.3)}{60}=\frac{\sin(A)}{76.45}\]Once you find A, you have two angles and you can subtract their sum from 180 to find C.

OpenStudy (anonymous):

on the other hand law of sines would be much faster =)

OpenStudy (stormfire1):

Maybe...I don't even remember the cosin theorem :P